# Phase transition on the fluctuation of the structure of random walk ranges

### v1

@@ 2. Preparations in $d\geq 5$
and $T^{1}$ (resp. $T^{-1}$) be the first global cut point after (resp. before) the origin (see the details in [4, Lemma 1.1]). Note that $T^{1}$ is the same distribution as that of $T^{-1}$ and
@@
then it suffices to show the result for $T^{1}$. In the equation above, we used the notation of [4], where $\omega$ is a variable in the state space that represents the sequence of steps taken by the random walk. $\theta^{n}\omega$ is the translation of the walk by $n$ time steps; namely, the new walk $\hat{S}$ given by $\theta^{n}\omega$ satisfies $\hat{S}_{i}=S_{i-n}$.
We now fix a parameter $J$, whose value will be set sufficiently large later. We then define $L=n/(3J)$ and set $k_{j}:=Lj$ for $0\leq j\leq 3J$.
If we let
@@
$a_{2}:=\mathbb{P}(\cap_{j=0}^{3J-1}\{N([1,n])=0,S(-\infty,k_{j}-1]\cap S[k_{j+1},\infty)=\emptyset\}),$
@@
(2.4) $a_{1}\leq C(J)n^{-\frac{d}{2}+2}.$
Note that if $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$, $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$ and $N([1,n])=0$, then $N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0$. To see the reason for this, note that $N([1,n])=0$ means that no point in the interval $k_{3j+1},k_{3j+2}-1]$ can be a cut point; thus, for $t\in[k_{3j+1},k_{3j+2}-1]$, we must have that $S(-\infty,t-1]\cap S[t,\infty)\neq\emptyset$. However, since $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$ and $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$, $S[k_{3j},t-1]\cap S[t,k_{3j+3}-1]\neq\emptyset$ holds instead. Hence,
@@
$A_{1}^{j,\alpha}=$ $\Big\{S[k_{3j},\,k_{3j+1}-1]\cap S[k_{3j+1}+n^{\alpha},k_{3j+3}-1]\neq\emptyset\Big\}$
$\cup\Big\{S[k_{3j},\,k_{3j+2}-n^{\alpha}]\cap S[k_{3j+2},\,k_{3j+3}-1]\neq\emptyset\Big\}.$
@@
Note that
$\mathbb{P}(A_{1}^{j,\alpha})\leq C(n^{\alpha})^{-\frac{d}{2}+2}.$
Now, we can split $k_{3j+1},k_{3j+2}-1]$” like in (2.3). That is, let $L^{\prime}=n^{\alpha}$, $J^{\prime}=\frac{k_{3j+2}-k_{3j+1}}{2L^{\prime}}$, $k^{\prime}_{\ell}=\ell L^{\prime}$ for $1\leq\ell\leq 2J^{\prime}$. Then, we have
@@
$\quad\leq\bigg\{\sum_{\ell=1}^{2J-1}\mathbb{P}\Big(S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{\ell}-1]\cap S[k_{3j+1}+k^{\prime}_{\ell+1},\,k_{3j+2}-1]\neq\emptyset\Big)\bigg\}(:=\hat{a}_{1})$
$\qquad+\mathbb{P}\Big(\text{for all }1\leq\ell\leq 2J^{\prime},\,S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{\ell-1}]\cap S[k_{3j+1}+k^{\prime}_{\ell},\,k_{3j+2}-1]=\emptyset,\,$
@@
$\hat{a}_{2}\leq$ $\mathbb{P}\Big(\text{for all }0\leq\ell\leq 2J^{\prime}+1,\,S(k_{3j},\,k_{3j+1}+k^{\prime}_{\ell-1}]\cap S[k_{3j+1}+k^{\prime}_{\ell},\,k_{3j+3})=\emptyset,\,$
@@
$\hat{a}_{1}\leq C\,n^{1-\alpha}\,(n^{\alpha})^{-\frac{d}{2}+2},\quad\hat{a}_{2}\leq C\,n^{-\alpha}.$
Then if we choose $\alpha$ close to $1$, for some $\delta>0$ (it should not depend on $J$),
$\hat{a}_{1}+\hat{a}_{2}\ll n^{-\delta}$
@@
$a_{2}\leq C(J)(n^{-\delta})^{J}.$
Therefore, if we pick sufficiently large $J$, we obtain $a_{2}\leq Cn^{-\frac{d}{2}+2}$ and in addition the desired result follows by combining (2.3) and (2.4). ∎
@@
$D_{1}:=\{\ X[n,2n]\geq c_{d}n\},$ $D_{2}:=\{{\rm Cap}(S[2n-sn,2n])\geq c^{\prime}_{d}sn\},$
$D_{3}:=\{\max_{2n-sn\leq i\leq 2n}\|S_{i}-S_{n}\|\leq C_{1}n^{1/2}\},$ $D_{4}:=\{\|S_{sn}-S_{n}\|\leq C_{1}n^{1/2}\}.$
Here, $c_{d}$ and $c_{d}^{\prime}$ are some dimension-dependent constants, (by the law of large numbers for the capacity of the random walk range, we choose $c_{d}^{\prime}$ to satisfy ${\rm Cap}(S[0,n])\sim 2c^{\prime}_{d}n$). Recall the (slightly modified) definition of $T^{1}$ as the first cut point after time $n$. By [7, Lemma 2.2], we know that there is $c_{d}>0$ such that $\mathbb{P}(X^{2}_{(-\infty,\infty)}[T^{1},n+T^{1}]\geq c_{d}n)=1-o(1)$. Noting that $X^{2}_{(-\infty,\infty)}[a,b]$ will be a lower bound for $X[a,b]$ for any times $a$ and $b$ and that $X^{2}_{(-\infty,\infty)}[a,b]$ is monotonically increasing as we increase the interval $[a,b]$, we obtain
@@
$\geq$ $\mathbb{P}(X^{2}_{(-\infty,\infty)}(0,n-f(n))\geq c_{d}n/2|0\text{ is cut point})-\mathbb{P}(T^{1}\geq n+f(n))=1-o(1)$
@@
On the event $D_{1}\cap D_{2}$, we see that we have $\max_{x\in A}\|x-z\|\leq 2C_{1}n^{1/2}$. We now reverse the random walk $S_{0},\ldots,S_{sn}$, to create a new random walk $\tilde{S}_{0},\ldots,\tilde{S}_{sn}$, defined as,
@@
$\geq\sum_{x\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=x)\mathbb{P}^{x}(T_{A}=\infty)$
$\geq\min_{a\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=a)\sum_{x\in A}\mathbb{P}^{x}(T_{A}=\infty)$
@@
On the event $D_{2}$, we assumed that the capacity of $A$ is greater than $sn$. Thus, conditionally on the intersection of the events $D_{2}$, $D_{3}$, and $D_{4}$, we can derive a lower bound on the probability of intersection between $\tilde{S}$ (which is also $S[1,sn]$) and A (which is S[2nsn,2n]S[2n-sn,2n]) as
@@
$X[n,2n]\geq c_{d}n,$
$X_{n}\geq 0,$
$X_{2n}\leq 2sn.$
The first line comes from the event $D_{1}$ and the second line is trivial. To see the last line (for the graph distance), notice that if we let $S_{t_{1}}$ and $S_{t_{2}}$ be the time indices of the intersection $\{S_{1},\ldots,S_{sn}\}$ with $\{S_{2n-sn},\ldots,S_{2n}\}$ with $t_{1}\in[0,sn]$ and $t_{2}\in[2n-sn,2n]$, the shortest geodesic must be smaller than $t_{1}+(2n-t_{2})$. Namely, shortest geodesic must be shorter than the path $S_{0},\ldots,S_{t_{1}},S_{t_{2}+1},\ldots,S_{2n}$. Now, $t_{1}$ and $2n-t_{2}$ are both less than $sn$. The same argument applies for the number of cut points. To see how the proof holds for the effective resistance, notice that since the effective resistance decreases upon adding edges, we must have that $X_{2n}$ is less than the effective resistance of the graph consisting only of the points $\{S_{1},\ldots,S_{t_{1}}\}\cup\{S_{t_{2}},\ldots,S_{2n}\}$. The effective resistance of this graph is less than $t_{1}+(2n-t_{2})$, the number of points along this graph.
Thus, we have that $E_{n}=X_{n}+X[n+1,2n]-X_{2n}\geq(c_{d}-2s)n$. Therefore,
$\mathbb{P}(E_{n}\geq 2(c_{d}-s)n)\geq\mathbb{P}(\{S[1,sn]\cap S[2n-sn,2n]\neq\emptyset\}\cap(\cap_{i=1}^{4}A_{i}))\geq c(s)n^{-d/2+2}.$
@@
We have the following estimates for the error terms $E_{n}$.
@@
As a consequence of the first and last inequality above, we have that $\mathrm{Var}(E_{n})\gtrsim n$.
Let $\Omega_{n}$ be the event that $\mathbb{E}[E_{n}]-E_{n}\geq 0$, then, we have that
@@
Since $E_{n}\geq 0$, we have that $\mathbb{E}[E_{n}]-E_{n}\lesssim\log n$. This gives us the final estimate we desire. ∎
@@ 3.1. Proof of variance
In this subsection, we will give the estimate of the variance.
@@
Using (2.1) and that $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq C(\log n)^{M}$ for $d\geq 8$ and $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq n^{1/4}$ for $d=7$, we find
@@
Namely, $N_{1}$ denotes the number of events in which $Y_{i}=1$ occurs exactly once and $N_{2}$ the number of events in which $Y_{i}=1$ appears exactly twice in a row. $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge n})_{l=1}^{\infty}$. For large $M$, there exists $c>0$ such that
@@
Eventually, for $|a-\mathbb{E}[N_{1}]|\leq M\sqrt{n}$, $|b-\mathbb{E}[N_{2}]|\leq M\sqrt{n}$,
@@
It holds that $\alpha_{d}>0$.
@@
and $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge(n/2)})_{l=1}^{\infty}$. By (3.1), there exists $c_{1}>0$ such that
@@
Since $\mathbb{E}X_{n}/n\to c_{0}$ by Remark 2.3, we have $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}\geq c_{3}$ for some $c_{3}>0$. Indeed, if $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}<c_{3}$ and set $a_{m}=\mathbb{E}X_{m^{2}}$, then $(a_{m}-a_{m-1})/m\leq c_{3}$ for large $m$ and hence $a_{m}\leq c_{0}m^{2}/2$ for large $m$ if we pick small $c_{3}$. We now pick a subsequence $n$ such that $(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})\geq c_{3}\sqrt{n}$ and choose $\epsilon<c_{3}/2$. Then,

### v2

@@ 2. Preparations in $d\geq 5$
and $T^{1}$ (resp. $T^{-1}$) be the first global cut point after (resp. before) the origin (see the details in [4, Lemma 1.1]). Note that $T^{1}$ has the same distribution (up to a sign change) as $T^{-1}$ and
@@
then it suffices to show the result for $T^{1}$. In the equation above, we used the notation of [4], where $\omega$ is a variable in the state space that represents the sequence of steps taken by the random walk. $\theta^{n}\omega$ is the translation of the walk by $n$ time steps; namely, the new walk $\hat{S}$ given by $\theta^{n}\omega$ satisfies $\hat{S}_{i}=S_{i-l}$.
We now fix a parameter $J$, whose value will be set sufficiently large later. We then define $L=l/(3J)$ and set $k_{j}:=Lj$ for $0\leq j\leq 3J$. If we let
@@
$a_{2}:=\mathbb{P}(\cap_{j=0}^{3J-1}\{N([1,l])=0,S(-\infty,k_{j}-1]\cap S[k_{j+1},\infty)=\emptyset\}),$
@@
(2.4) $a_{1}\leq C(J)l^{-\frac{d}{2}+2}.$
Note that if $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$, $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$ and $N([1,l])=0$, then $N([k_{3j+1},k_{3j+2}-1],[k_{3j},k_{3j+3}-1])=0$. To see the reason for this, note that $N([1,l])=0$ means that no point in the interval $k_{3j+1},k_{3j+2}-1]$ can be a cut point; thus, for $t\in[k_{3j+1},k_{3j+2}-1]$, we must have that $S(-\infty,t-1]\cap S[t,\infty)\neq\emptyset$. However, since $S(-\infty,k_{3j}-1]\cap S[k_{3j+1},\infty)=\emptyset$ and $S(-\infty,k_{3j+2}-1]\cap S[k_{3j+3},\infty)=\emptyset$, $S[k_{3j},t-1]\cap S[t,k_{3j+3}-1]\neq\emptyset$ holds instead. Hence,
@@
$A_{1}^{j,\alpha}=$ $\Big\{S[k_{3j},\,k_{3j+1}-1]\cap S[k_{3j+1}+l^{\alpha},k_{3j+3}-1]\neq\emptyset\Big\}$
$\cup\Big\{S[k_{3j},\,k_{3j+2}-l^{\alpha}]\cap S[k_{3j+2},\,k_{3j+3}-1]\neq\emptyset\Big\}.$
@@
By the same proof as that of (2.4),
$\mathbb{P}(A_{1}^{j,\alpha})\leq C(l^{\alpha})^{-\frac{d}{2}+2}.$
Now, we can split $k_{3j+1},k_{3j+2}-1]$” like in (2.3). That is, let $L^{\prime}=l^{\alpha}$, $J^{\prime}=\frac{k_{3j+2}-k_{3j+1}}{2L^{\prime}}$, $k^{\prime}_{u}=uL^{\prime}$ for $1\leq u\leq 2J^{\prime}$. Then, we have
@@
$\quad\leq\bigg\{\sum_{u=1}^{2J-1}\mathbb{P}\Big(S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{u}-1]\cap S[k_{3j+1}+k^{\prime}_{u+1},\,k_{3j+2}-1]\neq\emptyset\Big)\bigg\}(:=\hat{a}_{1})$
$\qquad+\mathbb{P}\Big(\text{for all }1\leq u\leq 2J^{\prime},\,S[k_{3j+1},\,k_{3j+1}+k^{\prime}_{u-1}]\cap S[k_{3j+1}+k^{\prime}_{u},\,k_{3j+2}-1]=\emptyset,\,$
@@
$\hat{a}_{2}\leq$ $\mathbb{P}\Big(\text{for all }0\leq u\leq 2J^{\prime}+1,\,S(k_{3j},\,k_{3j+1}+k^{\prime}_{u-1}]\cap S[k_{3j+1}+k^{\prime}_{u},\,k_{3j+3})=\emptyset,\,$
@@
$\hat{a}_{1}\leq C\,l^{1-\alpha}\,(n^{\alpha})^{-\frac{d}{2}+2},\quad\hat{a}_{2}\leq C\,l^{-\alpha}.$
Then, if we choose $\alpha$ close to $1$, for some $\delta>0$ (note that the obtained value of $\delta$ does not depend on $J$),
$\hat{a}_{1}+\hat{a}_{2}\ll l^{-\delta}$
@@
$a_{2}\leq C(J)(l^{-\delta})^{J}.$
Therefore, if we pick sufficiently large $J$, we obtain $a_{2}\leq Cl^{-\frac{d}{2}+2}$ and, in addition, the desired result, equation (2.1), follows by combining (2.3) and (2.4). ∎
@@
$D_{1}:=\{\ X[n,2n]\geq c_{d}n/2\},$ $D_{2}:=\{{\rm Cap}(S[2n-sn,2n])\geq c^{\prime}_{d}sn\},$
$D_{3}:=\{\max_{2n-sn\leq i\leq 2n}\|S_{i}-S_{n}\|\leq C_{1}n^{1/2}\},$ $D_{4}:=\{\|S_{sn}-S_{n}\|\leq C_{1}n^{1/2}\},$
(see the definition of the capacity in [12] et al.). Here, $c_{d}$ and $c_{d}^{\prime}$ are some dimension-dependent constants (by the law of large numbers for the capacity of the random walk range, we choose $c_{d}^{\prime}$ to satisfy ${\rm Cap}(S[0,n])\sim 2c^{\prime}_{d}n$). In addition, we choose $0<s<c_{d}/8$. Recall the (slightly modified) definition of $T^{1}$ as the first cut point after time $n$. By [7, Lemma 2.2], we know that there is $c_{d}>0$ such that $\mathbb{P}(X^{2}_{(-\infty,\infty)}[T^{1},n+T^{1}]\geq c_{d}n)=1-o(1)$. Noting that $X^{2}_{(-\infty,\infty)}[a,b]$ will be a lower bound for $X[a,b]$ for any times $a$ and $b$ and that $X^{2}_{(-\infty,\infty)}[a,b]$ is monotonically increasing as we increase the interval $[a,b]$, we obtain
@@
$\geq$ $\mathbb{P}(X^{2}_{(-\infty,\infty)}[0,n-f(n)]\geq c_{d}n/2|0\text{ is cut point})-\mathbb{P}(T^{1}\geq n+f(n))=1-o(1)$
@@
On the event $D_{3}\cap D_{4}$, we see that we have $\max_{x\in A}\|x-z\|\leq 2C_{1}n^{1/2}$. We now reverse the random walk $S_{0},\ldots,S_{sn}$, to create a new random walk $\tilde{S}_{0},\ldots,\tilde{S}_{sn}$, defined as,
@@
$\geq\sum_{x\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=x)\mathbb{P}^{x}(T_{A}=\infty)\geq\min_{a\in A}\sum_{i=1}^{sn}\mathbb{P}^{z}(\tilde{S}_{i}=a)\sum_{x\in A}\mathbb{P}^{x}(T_{A}=\infty)$
@@
On the event $D_{2}$, we assumed that the capacity of $A$ is greater than $c_{d}^{\prime}sn$. Thus, conditionally on the intersection of the events $D_{2}$, $D_{3}$, and $D_{4}$, we can derive a lower bound on the probability of intersection between $\tilde{S}$ (which is also $S[1,sn]$) and $A$ (which is S[2nsn,2n]S[2n-sn,2n]) as
@@
$X[n,2n]\geq c_{d}n/2,\quad X_{n}\geq 0,\quad X_{2n}\leq 2sn.$
The first line comes from the event $D_{1}$, and the second line is trivial. To see the last line (for the graph distance), notice that if we let $S_{t_{1}}$ and $S_{t_{2}}$ be the points of the intersection $\{S_{1},\ldots,S_{sn}\}$ and $\{S_{2n-sn},\ldots,S_{2n}\}$ with $t_{1}\in[0,sn]$ and $t_{2}\in[2n-sn,2n]$, the shortest geodesic must be smaller than $t_{1}+(2n-t_{2})$. Namely, the shortest geodesic must be shorter than the path $S_{0},\ldots,S_{t_{1}},S_{t_{2}+1},\ldots,S_{2n}$. Now, $t_{1}$ and $2n-t_{2}$ are both less than $sn$. The same argument applies to the number of cut points.
Thus, we have that $E_{n}=X_{n}+X[n+1,2n]-X_{2n}\geq(c_{d}/2-2s)n$. Therefore,
$\mathbb{P}(E_{n}\geq 2(c_{d}/2-s)n)\geq\mathbb{P}(\{S[1,sn]\cap S[2n-sn,2n]\neq\emptyset\}\cap(\cap_{i=1}^{4}A_{i}))\geq c(s)n^{-d/2+2}.$
@@
We have the following estimates for the error terms $E_{n}$:
@@
As a consequence of the first and last inequality above, $\mathrm{Var}(E_{n})\gtrsim n$.
Let $\Omega_{n}$ be the event that $\mathbb{E}[E_{n}]-E_{n}\geq 0$, then,
@@
Since $E_{n}\geq 0$, we obtain that $\mathbb{E}[E_{n}]-E_{n}\lesssim\log n$. This gives us the final estimate we desire. ∎
@@ 3.1. Proof of variance
In this subsection, we will give an estimate for the variance.
@@
Using (2.1) and that $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq C(\log n)^{M}$ with $M<\infty$ for $d\geq 8$ and $\max\left(\|E_{n}\|_{2},\mathbb{E}[E_{n}]\right)\leq n^{1/4}$ for $d=7$, we find
@@
Namely, $N_{1}$ denotes the number of events in which $Y_{i}=1$ occurs exactly once, and $N_{2}$ denotes the number of events in which $Y_{i}=1$ appears exactly twice in a row. $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge n})_{l=1}^{\infty}$.
Now, we state our desired conclusions. For large $M$, there exists $c>0$ such that
@@
As a consequence, for $|a-\mathbb{E}[N_{1}]|\leq M\sqrt{n}$, $|b-\mathbb{E}[N_{2}]|\leq M\sqrt{n}$,
@@
It holds that $\alpha_{d}>0$ for $d\geq 7$.
@@
and $\mathbb{P}^{n}$ denotes the probability for $(Y_{T_{l}\wedge(n/2)})_{l=1}^{\infty}$. $N_{1}$ counts the number of backtracks, while $N_{2}$ counts the number of double backtracks. We remark that changing a single backtrack to a double backtrack will not change the graph distance of the walk. However, the act of changing all single backtracks to double backtracks and vice versa would substantially change the length of the random walk. The existence of a procedure to change the volume without changing the graph distance essentially allows us to argue that the variance of the graph distance must be non-trivial.
By (3.1), there exists $c_{1}>0$ such that
@@
Since $\mathbb{E}X_{n}/n\to c_{0}$ by Remark 2.3, we have $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}\geq c_{3}$ for some $c_{3}>0$. Indeed, if $\limsup_{n\to\infty}(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})/\sqrt{n}<c_{3}$ and we set $a_{m}=\mathbb{E}X_{m^{2}}$, then $(a_{m}-a_{m-1})/m\leq c_{3}$ for large $m$; hence, $a_{m}\leq c_{0}m^{2}/2$ for large $m$ if we pick $c_{3}$ were chosen small enough. We now pick a subsequence $n$ such that $(\mathbb{E}X_{n}-\mathbb{E}X_{n-\sqrt{n}})\geq c_{3}\sqrt{n}$ and choose $\epsilon<c_{3}/2$. Then,